# Efficient Forkless Blockchain Databases

### v2

@@ Abstract.
In this paper, we propose a forkless blockchain database, showing a 100x improvement in storage and a 10x improvement in throughput compared to the geth-based Fantom Blockchain client.
@@ 1. Introduction
Early proof-of-work blockchains, such as Bitcoin and Ethereum, are hampered by limited performance, i.e., achieving less than 30 transactions per second (TPS) (Defiant, 2024). These blockchains have forking semantics; when the ledger’s consistency cannot be temporarily enforced, multiple forks are maintained and processed using different versions of the account’s state. Eventually, a single fork is globally selected based on a heuristic, such as the longest-chain rule (Nakamoto, 2009).
In an effort to improve performance, modern blockchains have adopted forkless proof-of-stake consensus protocols. In these blockchains, the consensus protocol ensures that a consistent set of blocks is produced among the nodes, thereby avoiding forking. The vast majority of modern blockchains incorporate proof-of-stake consensus protocols (Ge et al., 2022; Buchman et al., 2022; Spiegelman et al., 2022; Castro and Liskov, 1999; Babel et al., 2025; Baird and Luykx, 2020; Yakovenko, 2024) to achieve a throughput of thousands of transactions per second (Baird et al., 2019).
However, forkless blockchains continue to use legacy databases that were initially designed for forking blockchains only. Recent profiling experiences (Baird et al., 2019) report that transaction execution in forkless blockchains is now primarily dominated by storage access, which consumes up to 75% of the block processing. Moreover, they require nodes to store terabytes of data [^1] Storing the entire history of the Ethereum mainnet currently requires 24 TB of disk space, according to https://etherscan.io/chartsync/chainarchive.
@@
However, the MPTs ability to track multiple versions of a block state with multiple roots is costly as it requires pruning (i.e., disposing of worldstates that are no longer needed) and copying of data. And yet, for forkless blockchains, this behaviour is superfluous. As a result of this one-size-fits-all solution, forkless blockchains experience limited throughput and require nodes with terabytes of expensive storage. We demonstrate this phenomenon below in Example 1.1.
Example 1.1.
@@
Note that the highlighted path for a forkless blockchain is redundant, as it is not needed for the new root $x_{2}$. Retaining the data causes bloat and thus will eventually be pruned (Ethereum, 2025), which typically requires putting the node offline. Moreover, as shown by the dotted arrows, numerous copies of keys along the path need to be copied to create a path from $x_{2}$ to the new value $v^{\prime}_{4}$. Overall, considering an MPT typically holds hundreds of millions of values, the highlighted inefficiencies lead to considerable computational overhead and redundant bloat if previous versions are not needed.
@@
The key aspect of our database design is that the LiveDB uses a mutable data structure that allows data to be overwritten, reflecting the constrained versioning occurring in forkless blockchains. Overwriting allows us to reflect the state of the most recent execution without keeping stale states in the database (right MPT in Figure 1). Thus, we perform pruning at no additional cost with LiveDB. We refer to this method as intrinsic pruning that is inherently performed as a consequence of the chosen data structure. Moreover, the LiveDB stores key-value maps as densely as possible. A dense representation of the key-value maps results in smaller disk footprints and faster accesses. This design is in stark contrast to the MPT implementations that use key-value stores (e.g., LevelDB (Inc., 2023a), RocksDB (Inc., 2023b)) and SQL databases.
@@
We have implemented LiveDB and ArchiveDB in the Fantom Blockchain and evaluated their utility on real world transactions in a public testnet. Our evaluation shows that LiveDB reduces the storage footprint [^2] Note that even smaller storage decreases imply significant storage reductions linear in the number of nodes in the network because each node replicates the data in a classical L1 blockchain. on the Fantom blockchain by approximately $100\times$ while increasing throughput $10\times$ compared to using the MPT-based Ethereum geth client (Ethereum, 2023).
@@ 2. StateDB Specialization
Nodes in a blockchain have different roles. A node is assigned a role by configuring the software, called the client, that operates on each node. Validator roles are tasked with producing a new block in the blockchain. To achieve this, validators run a proof-of-stake consensus protocol (Bano et al., 2019), which adds new transactions to the ledger and consistently orders them into blocks. The block processor of the validator executes the transactions using a virtual machine and a blockchain database (StateDB).
@@
A typical block processing design is illustrated in Figure 2(a), utilizing an unspecialized StateDB design employed in chains such as Ethereum (Community, 2020) and Fantom (Choi et al., 2018). Here, the consensus component generates new blocks, enabling the block processor, which includes the virtual machine and the StateDB, to advance to the next state. In this block processing architecture, a single StateDB is used to interact with both the Virtual Machine (VM) and external RPC calls for historical queries, such as querying the balance of an account in a specific block. A standard implementation for a StateDB is based on Merkle-Patricia Trie (Merkle, 1989) that supports both archival and live data storage. However, the configuration in Figure 2(a) results in excessive storage overhead for validator and observer nodes that do not need historical states.
@@
In Figure 2(b), validators and observers only have a LiveDB but no ArchiveDB. The VM will access and update the LiveDB when processing the next block. The LiveDB maintains a single state for its accounts and has no multiple worldstates. In Figure 2(c), the ArchiveDB is used to process historical RPC requests and is read-only by the RPC server. When a new block arrives, block processing generates new states using the LiveDB and writes the updated accounts for a block into the ArchiveDB.
@@ 3. LiveDB with Intrinsic-Pruning
The LiveDB design has two design principles. The first is a normalization (Abiteboul et al., 1995) step where the attributes of accounts are not directly fetched. Instead, an account address is mapped to a record number, and with the record number, the account attribute is retrieved. For smart contract storage, we use an account address and storage key mapping that maps to the record number of the storage value. This allows for the direct storage of account attributes and their corresponding values. Instead of using a B-Tree or other ordered data structures with ${\mathcal{O}}(\log n)$, we use an efficient I/O hashmap (Henderson, [n. d.]) for the normalization mapping account addresses/storage keys to record numbers.
@@
There are hashing strategies, such as linear hashing (Henderson, [n. d.]), that are I/O-friendly, allowing for efficient hash-table growth via splitting without requiring a complete reconstruction of the hash table. This is achieved by expanding the bits of the index calculation depending on the hash table size. The hash table stores the keys and their corresponding record numbers, allowing for the retrieval of account records or storage values. If the stored key in the hashtable is not identical to the searched key, the subsequent entries are searched until the key is found. The hash table stores the account address/storage key and the record number of the attributes/storage value as shown in Figure 4(b). The key is needed so that, in the event of a collision caused by the hash function, we can find a neighboring slot for storing the key and its ordinal number.
@@ 3.1. Hash Calculations
For the file-mapped arrays of attributes and smart contract storage values, an underlying balanced binary tree is constructed for hash calculations using the storage/attribute values in a binary file. We refer to the balanced binary trees as the hash trees of the attributes and storage values. They are also illustrated in Figure 3. We can construct balanced binary trees using ideas from the heap data structure, as employed in heap sorts and other algorithms (Boothroyd, 1964; Cormen et al., 2009), by utilizing arrays and index calculations to represent parent-child relations. Hence, hash trees can also be implemented as file-mapped arrays. Additionally, the balanced hash tree can be incrementally extended to accommodate new attribute/storage value entries without compromising the integrity of large portions of the hashes.
@@
The advantage of a lazy update is that for a set of changed attributes/storage values, the hash computation is performed in one go by visiting a node in the hashtree at most once. This improves the amortized complexity and is achieved by using a reverse breadth-first traversal, which can be fully parallelized using the ideas of prefix-sum (Ladner and Fischer, 1980).
@@ 5. Implementation
  • HashTree: A structure that operates on the PagePool to compute Merkle proofs(Merkle, 1989) from the pages.
@@ 6. Evaluation
We conducted a series of experiments to evaluate transaction throughput and the disk storage space required by various implementations of our technique. For this purpose, we adapted the tool for blockchain recording and replaying (Kim et al., 2021) to record blockchain data available on the Fantom mainnet. This dataset comprised 50 million blocks, which were stored locally on an SSD NVMe disk. The tool was subsequently used to reprocess these blocks through selected components of the Fantom client, thereby reconstructing the blockchain. Specifically, we modified the block processing module, which encompasses the Ethereum Virtual Machine (EVM) and the database, to isolate the influence of other components, such as the network, consensus protocol, and transaction pool. This allowed us to precisely measure the number of transactions processed within a sampled time frame and the corresponding disk space consumed.
@@ 6.1. Experimental Setup
We employed the tool (Kim et al., 2021), which is designed to operate in two distinct phases. In the first phase, the tool synchronized an existing blockchain and recorded the necessary input data to enable the re-execution of all transactions in each block. This process was applied to the Fantom mainnet, a blockchain that has been operational since 2018. Notably, this recording phase was not included in the performance measurements. In the second phase, the tool leveraged the recorded data to replay the entire blockchain at an accelerated rate, wherein the software and hardware imposed the only constraints. This approach provides insights into key performance metrics, including the number of transactions processed per second and the volume of blockchain gas consumed over a given period.
@@ 7. Related Work
The Merkle Patricia Trie (MPT) is a fundamental data structure in blockchain systems, combining the advantages of Merkle trees and Patricia tries. Merkle trees, introduced by Merkle (Merkle, 1989), provide a cryptographic method to ensure data integrity by hashing values of nodes in the trie. Patricia, as proposed by Morrison (Morrison, 1968), offers an efficient method for compressing sparse data by compressing common paths of keys into shared sequences. Read and write amplification in tries has been a significant research focus. Raju (Raju, 2018) introduced the Merkelized LSM, integrating Merkle and log-structured merge tree data structures. Ponnapalli (Ponnapalli, 2019) proposed a distributed Merkle Patricia trie across multiple nodes. Extensive research has optimized Merkle tree structures (Dahlberg et al., 2016; Haider et al., 2018; Östersjö et al., 2016). Despite these advancements, our experiments suggest that the performance benefits are limited, as trie lookup is significantly impacted by database latency.
@@
Cache-Tries, introduced by Prokopec (Prokopec, 2018), present a design for concurrent hash tries that ensures constant-time operations using a quiescently consistent cache. This approach extends the cache-trie with single-linked lists, each holding pointers to nodes at one level of the trie, sorted from the deeper levels to the root. This cache enables quick lookup and insertion by inspecting the tree from the bottom up, starting from the deepest level, or reverting to a slower top-down search from the root.
@@
Efficient privacy-preserving storage, as proposed by BlockStorage (Jia et al., 2023), introduces a blockchain storage method tailored for IoT environments. This method leverages federated extreme learning machines to classify and store hot blocks efficiently while preserving user privacy. This approach enhances data query speed and ensures that sensitive information remains protected from unauthorized access.
@@
Feng et al. (Feng et al., 2023) propose a space-efficient storage structure for blockchain transactions that supports secure verification. Their approach minimizes the storage footprint of transaction data while ensuring reliable verification processes. By employing advanced data structures and algorithms, their method significantly reduces the required storage space without compromising blockchain security and integrity.
@@
Li et al. (Li et al., 2024) propose a replacement to the MPT data structure that, like our approach, avoids the use of an off-the-shelf key/value store, resulting in ${\mathcal{O}}(\log_{2}n)$ lookup complexity. They instead propose a solution that utilizes a Layered Versioned Multipoint Trie data structure, resulting in ${\mathcal{O}}(\log n)$ lookup complexity. Prior to (Li et al., 2024), there was work on similar commit data structures with better performance than MPT (Raju et al., 2018).
@@
A substantial body of work has explored concurrency to address performance limitations of smart contracts (Dickerson, 2017; Zhang et al., 2018; Anjana et al., 2019; Shrey, 2019; Sarrar, 2019; Saraph, 2019). Miners construct an execution graph with forks and joins, which is replicated by other network nodes. This strategy removes database access from the critical path, thereby enhancing overall performance; however, it does not mitigate the latency associated with retrieving individual state values from the database.
@@
Research has also focused on developing faster key-value databases. Faster (Chandramouli et al., 2018) is a concurrent database that spans both hard drives and main memory with a log. Song et al. (Song et al., 2018) introduced bulk get methods atop RocksDB, optimizing SSD usage. Papagiannis et al. (Papagiannis et al., 2018) reported a 4–6 times speedup over RocksDB with their memory-mapped key-value store. Ouaknine et al. (Ouaknine et al., 2017) showed substantial performance improvements with a highly-tuned RocksDB configuration on SSDs. Wu et al. (Wu et al., 2015) proposed an LSM-trie algorithm, achieving a tenfold performance enhancement with LevelDB and RocksDB, with no performance degradation as data size increased. Lepers et al. (Lepers et al., 2019) introduced a database optimized for modern SSDs, while Agarwal et al. (Agarwal et al., 2015) suggested querying compressed data to save disk space without performance loss.
@@
The Ethereum community is exploring sharding to distribute computational load and reduce the size of the state trie (Community, 2020). Sharding segments nodes into clusters by account, allowing shards to operate independently. Recently, Ethereum has also focused on a slower L1 chain, offloading scalability to L2 chains. The Geth developers have proposed state snapshots by enabling a separate index [^5] https://github.com/ethereum/go-ethereum/pull/20152. Support for live pruning in MPT, based on the Serial index, is documented by the Klaytn Foundation  [^6] https://docs.klaytn.foundation/docs/build/tools/indexers/, which involves the selective removal of obsolete data to optimize storage and performance. The Path-Based scheme, discussed in Ethereum research forums [^7] https://ethresear.ch/t/ethereum-path-based-storage-model-and-newly-inline-state-prune/14932 and implemented in the Geth client [^8] https://github.com/ethereum/go-ethereum/issues/25390, introduces a mutable MPT scheme for live databases, enhancing the trie with embedded pruning capability.
@@
One way to ensure that we keep the storage overheads and access time low for validators and observers is to perform live pruning (Ethereum, 2025) in the MPT data structure of the StateDB, as there are no forks and the evolution of blocks is linear. After a new version is introduced, the unused nodes of the old version are removed from the MPT so that only the latest worldstate of the blockchain is kept in the MPT. Live pruning is beneficial for validators and observers, as they only require the most recent state in the StateDB for a forkless chain. However, live pruning is a form of garbage collection that involves a node traversal within the MPT, collecting outdated nodes and eliminating them. In case of a multi-threaded client, locking must be implemented to avoid data races when live pruning is performed in parallel with new value updates, which complicates the implementation of the MPT and worsens performance. On the other hand, ArchiveDB does not require live pruning for forkless chains [^10] Only for forked chains, some forks could be removed to safe storage, applying the longest-chain rule.. The worldstates of all blocks must be retained in the MPT. However, for forkless chains, the evolution of the versions is linear.

### v3

@@ Abstract.
In this paper, we propose a forkless blockchain database that avoids costly operations associated with forking. Our improved database structure exhibits a 100x improvement in storage and a 10x improvement in throughput compared to the geth-based Fantom Blockchain client.
@@ 1. Introduction
Early proof-of-work blockchains, such as Bitcoin and Ethereum, are hampered by limited performance, i.e., achieving less than 30 transactions per second (TPS) (ethtps). These blockchains have forking semantics; when the ledger’s consistency cannot be temporarily enforced, multiple forks are maintained and processed using different versions of the account’s state. Eventually, a single fork is globally selected based on a heuristic, such as the longest-chain rule (Nakamoto2009BitcoinA).
In an effort to improve performance, modern blockchains have adopted forkless proof-of-stake consensus protocols. In these blockchains, the consensus protocol ensures that a consistent set of blocks is produced among the nodes, thereby avoiding forking. The vast majority of modern blockchains incorporate proof-of-stake consensus protocols (10.1155/2022/2812526; tendermint; bullshark; pbft; mys; BairdL20; solana) to achieve a throughput of thousands of transactions per second (abs-1910-11143).
However, forkless blockchains continue to use legacy databases that were initially designed for forking blockchains only. Recent profiling experiences (abs-1910-11143) report that transaction execution in forkless blockchains is now primarily dominated by storage access, which consumes up to 75% of the block processing. Moreover, they require nodes to store terabytes of data [^1] Storing the entire history of the Ethereum mainnet currently requires 24 TB of disk space, according to https://etherscan.io/chartsync/chainarchive.
@@
However, the MPTs ability to track multiple versions of a block state with multiple roots is costly as it requires pruning (i.e., disposing of worldstates that are no longer needed) and copying of data. And yet, for forkless blockchains, this behavior is superfluous. As a result of this one-size-fits-all solution, forkless blockchains experience limited throughput and require nodes with terabytes of expensive storage. We demonstrate this phenomenon below in Example 1.1.
Example 0.
@@
Note that the highlighted path for a forkless blockchain is redundant, as it is not needed for the new root $x_{2}$. Retaining the data causes bloat and thus will eventually be pruned (livepruning), which typically requires putting the node offline. Moreover, as shown by the dotted arrows, numerous copies of keys along the path need to be copied to create a path from $x_{2}$ to the new value $v^{\prime}_{4}$. Overall, considering an MPT typically holds hundreds of millions of values, the highlighted inefficiencies lead to considerable computational overhead and redundant bloat if previous versions are not needed.
@@
The key aspect of our database design is that the LiveDB uses a mutable data structure that allows data to be overwritten, reflecting the constrained versioning occurring in forkless blockchains. Overwriting allows us to reflect the state of the most recent execution without keeping stale states in the database (right MPT in Figure 1). Thus, we perform pruning at no additional cost with LiveDB. We refer to this method as intrinsic pruning that is inherently performed as a consequence of the chosen data structure. Moreover, the LiveDB stores key-value maps as densely as possible. A dense representation of the key-value maps results in smaller disk footprints and faster accesses. This design is in stark contrast to the MPT implementations that use key-value stores (e.g., LevelDB (leveldb), RocksDB (rocksdb)) and SQL databases.
@@
We have implemented LiveDB and ArchiveDB in the Fantom Blockchain and evaluated their utility on real world transactions in a public testnet. Our evaluation shows that LiveDB reduces the storage footprint [^2] Note that even smaller storage decreases imply significant storage reductions linear in the number of nodes in the network because each node replicates the data in a classical L1 blockchain. on the Fantom blockchain by approximately $100\times$ while increasing throughput $10\times$ compared to using the MPT-based Ethereum geth client (geth).
@@ 2. StateDB Specialization
Nodes in a blockchain have different roles. A node is assigned a role by configuring the software, called the client, that operates on each node.
Validator roles are tasked with producing a new block in the blockchain. To achieve this, validators run a proof-of-stake consensus protocol (POS), which adds new transactions to the ledger and consistently orders them into blocks. The block processor of the validator executes the transactions using a virtual machine and a blockchain database (StateDB).
@@
A typical block processing design is illustrated in Figure 2(a), utilizing an unspecialized StateDB design employed in chains such as Ethereum and Fantom (fantom). Here, the consensus component generates new blocks, enabling the block processor, which includes the virtual machine and the StateDB, to advance to the next state. In this block processing architecture, a single StateDB is used to interact with both the Virtual Machine (VM) and external RPC calls for historical queries, such as querying the balance of an account in a specific block. A standard implementation for a StateDB is based on Merkle-Patricia Trie (merkle) that supports both archival and live data storage. However, the configuration in Figure 2(a) results in excessive storage overhead for validator and observer nodes that do not need historical states.
@@
In Figure 2(b), validators and observers only have a LiveDB but no ArchiveDB. The VM will access and update the LiveDB when processing the next block. The LiveDB maintains a single state for its accounts and has no multiple worldstates.
In Figure 2(c), the ArchiveDB is used to process historical RPC requests and is read-only by the RPC server. When a new block arrives, block processing generates new states using the LiveDB and writes the updated accounts for a block into the ArchiveDB.
@@ 3. LiveDB with Intrinsic-Pruning
The LiveDB design has two design principles. The first is a normalization (10.5555/551350) step where the attributes of accounts are not directly fetched. Instead, an account address is mapped to a record number, and with the record number, the account attribute is retrieved. For smart contract storage, we use an account address and storage key mapping that maps to the record number of the storage value. This allows for the direct storage of account attributes and their corresponding values. Instead of using a B-Tree or other ordered data structures with ${\mathcal{O}}(\log n)$, we use an efficient I/O hashmap (linear-hashing) for the normalization mapping account addresses/storage keys to record numbers.
@@
There are hashing strategies, such as linear hashing (linear-hashing), that are I/O-friendly, allowing for efficient hash-table growth via splitting without requiring a complete reconstruction of the hash table. This is achieved by expanding the bits of the index calculation depending on the hash table size. The hash table stores the keys and their corresponding record numbers, allowing for the retrieval of account records or storage values. If the stored key in the hashtable is not identical to the searched key, the subsequent entries are searched until the key is found. The hash table stores the account address/storage key and the record number of the attributes/storage value as shown in Figure 4(b). The key is needed so that, in the event of a collision caused by the hash function, we can find a neighboring slot for storing the key and its ordinal number.
@@ 3.1. Hash Calculations
For the file-mapped arrays of attributes and smart contract storage values, an underlying balanced binary tree is constructed for hash calculations using the storage/attribute values in a binary file. We refer to the balanced binary trees as the hash trees of the attributes and storage values. They are also illustrated in Figure 3. We can construct balanced binary trees using ideas from the heap data structure, as employed in heap sorts and other algorithms (10.1145/512274.512284; 10.5555/1614191), by utilizing arrays and index calculations to represent parent-child relations. Hence, hash trees can also be implemented as file-mapped arrays. Additionally, the balanced hash tree can be incrementally extended to accommodate new attribute/storage value entries without compromising the integrity of large portions of the hashes. Whenever a new attribute is inserted, it is appended at the end of the array, consequently creating a new branch of the hashing tree siting above this array. The array itself is always expanded by chunks to accommodate branching factor of the hashing tree, and the new chunk is created when the previous one is filled. This new chunk builds a new level of the hashing tree appended at the top of the tree, thus only this part of the tree must be recomputed.
@@
The advantage of a lazy update is that for a set of changed attributes/storage values, the hash computation is performed in one go by visiting a node in the hashtree at most once. This improves the amortized complexity and is achieved by using a reverse breadth-first traversal, which can be fully parallelized using the ideas of prefix-sum (10.1145/322217.322232).
@@ 4.1. Hash Calculations
The hashing mechanism allows for checking consistency of the ArchiveDB at any block height given ${\mathcal{O}}(1)$ access time complexity, while proving particular values of accounts and storage slots need a linear scan of the dataset.
@@ 5. Implementation
  • HashTree: A structure that operates on the PagePool to compute Merkle proofs(merkle) from the pages.
@@ 6. Evaluation
We conducted a series of experiments to evaluate transaction throughput and the disk storage space required by various implementations of our technique. For this purpose, we adapted the tool for blockchain recording and replaying (Kim2021off-the-chain) to record blockchain data available on the Fantom mainnet. This dataset comprised 50 million blocks, which were stored locally on an SSD NVMe disk. The tool was subsequently used to reprocess these blocks through selected components of the Fantom client, thereby reconstructing the blockchain. Specifically, we modified the block processing module, which encompasses the Ethereum Virtual Machine (EVM) and the database, to isolate the influence of other components, such as the network, consensus protocol, and transaction pool. This allowed us to precisely measure the number of transactions processed within a sampled time frame and the corresponding disk space consumed.
@@ 6.1. Experimental Setup
We employed the tool (Kim2021off-the-chain), which is designed to operate in two distinct phases. In the first phase, the tool synchronized an existing blockchain and recorded the necessary input data to enable the re-execution of all transactions in each block. This process was applied to the Fantom mainnet, a blockchain that has been operational since 2018. Notably, this recording phase was not included in the performance measurements. In the second phase, the tool leveraged the recorded data to replay the entire blockchain at an accelerated rate, wherein the software and hardware imposed the only constraints. This approach provides insights into key performance metrics, including the number of transactions processed per second and the volume of blockchain gas consumed over a given period.
@@ 7. Related Work
The Merkle Patricia Trie (MPT) is a fundamental data structure in blockchain systems, combining the advantages of Merkle trees and Patricia tries. Merkle trees, introduced by Merkle (merkle), provide a cryptographic method to ensure data integrity by hashing values of nodes in the trie. Patricia, as proposed by Morrison (Morrison68), offers an efficient method for compressing sparse data by compressing common paths of keys into shared sequences. Read and write amplification in tries has been a significant research focus. Raju (raju2018mlsm) introduced the Merkelized LSM, integrating Merkle and log-structured merge tree data structures. Ponnapalli (ponnapalli2019scalable) proposed a distributed Merkle Patricia trie across multiple nodes. Extensive research has optimized Merkle tree structures (haider2018compact; ostersjo2016sparse). Despite these advancements, our experiments suggest that the performance benefits are limited, as trie lookup is significantly impacted by database latency.
@@
Cache-Tries, introduced by Prokopec (hashtrie), present a design for concurrent hash tries that ensures constant-time operations using a quiescently consistent cache. This approach extends the cache-trie with single-linked lists, each holding pointers to nodes at one level of the trie, sorted from the deeper levels to the root. This cache enables quick lookup and insertion by inspecting the tree from the bottom up, starting from the deepest level, or reverting to a slower top-down search from the root.
@@
Efficient privacy-preserving storage, as proposed by Jia et al. (blockstorage), introduces a blockchain storage method tailored for IoT environments. This method leverages federated extreme learning machines to classify and store hot blocks efficiently while preserving user privacy. This approach enhances data query speed and ensures that sensitive information remains protected from unauthorized access.
@@
Feng et al. (Feng23) propose a space-efficient storage structure for blockchain transactions that supports secure verification. Their approach minimizes the storage footprint of transaction data while ensuring reliable verification processes. By employing advanced data structures and algorithms, their method significantly reduces the required storage space without compromising blockchain security and integrity.
@@
Li et al. (lvmt) propose a replacement to the MPT data structure that, like our approach, avoids the use of an off-the-shelf key/value store, resulting in ${\mathcal{O}}(\log_{2}n)$ lookup complexity. They instead propose a solution that utilizes a Layered Versioned Multipoint Trie data structure, resulting in ${\mathcal{O}}(\log n)$ lookup complexity.
Hang et al. (hang2024slimarchive) introduced a work resembling ours. They claim that the witness proof is often not necessary in many use-cases, and propose an archive that stores only account and storage data without the MPT hashes. This easily produces a compact archive that consumes only a fraction of the full archive disk space.
Zhang et al. (zhang2024cle) proposes a new indexing mechanism for MPT based on learned model, which is smaller in size and faster for querying.
@@
A substantial body of work has explored concurrency to address performance limitations of smart contracts (dickerson2017adding; anjana2019efficient; baheti2022dipetrans; saraph2019empirical). Miners construct an execution graph with forks and joins, which is replicated by other network nodes. This strategy removes database access from the critical path, thereby enhancing overall performance; however, it does not mitigate the latency associated with retrieving individual state values from the database.
@@
Research has also focused on developing faster key-value databases. Faster (chandramouli2018faster) is a concurrent database that spans both hard drives and main memory with a log. Song et al. (song2018multipath) introduced bulk get methods atop RocksDB, optimizing SSD usage. Papagiannis et al. (papagiannis2018efficient) reported a 4–6 times speedup over RocksDB with their memory-mapped key-value store. Ouaknine et al. (ouaknine2017optimization) showed substantial performance improvements with a highly-tuned RocksDB configuration on SSDs. Wu et al. (wu2015lsm) proposed an LSM-trie algorithm, achieving a tenfold performance enhancement with LevelDB and RocksDB, with no performance degradation as data size increased. Lepers et al. (lepers2019kvell) introduced a database optimized for modern SSDs, while Agarwal et al. (agarwal2015succinct) suggested querying compressed data to save disk space without performance loss.
@@
The Ethereum community is exploring sharding to distribute computational load and reduce the size of the state trie (yu2020survey). Sharding segments nodes into clusters by account, allowing shards to operate independently. Recently, Ethereum has also focused on a slower L1 chain, offloading scalability to L2 chains. Serial index, documented by the Klaytn Foundation  [^5] https://docs.klaytn.foundation/docs/build/tools/indexers/, involves the selective removal of obsolete data to optimize storage and performance. The Path-Based scheme, discussed in Ethereum research forums [^6] https://ethresear.ch/t/ethereum-path-based-storage-model-and-newly-inline-state-prune/14932 and implemented in the Geth client [^7] https://github.com/ethereum/go-ethereum/issues/25390, introduces a mutable MPT scheme for live databases, enhancing the trie with embedded pruning capability.
@@
One way to ensure that we keep the storage overheads and access time low for validators and observers is to perform live pruning (livepruning) in the MPT data structure of the StateDB, as there are no forks and the evolution of blocks is linear. After a new version is introduced, the unused nodes of the old version are removed from the MPT so that only the latest worldstate of the blockchain is kept in the MPT. Live pruning is beneficial for validators and observers, as they only require the most recent state in the StateDB for a forkless chain. However, live pruning is a form of garbage collection that involves a node traversal within the MPT, collecting outdated nodes and eliminating them. In case of a multi-threaded client, locking must be implemented to avoid data races when live pruning is performed in parallel with new value updates, which complicates the implementation of the MPT and worsens performance. On the other hand, ArchiveDB does not require live pruning for forkless chains [^9] Only for forked chains, some forks could be removed to safe storage, applying the longest-chain rule.. The worldstates of all blocks must be retained in the MPT. However, for forkless chains, the evolution of the versions is linear.